OpenStudy (anonymous):

f(x) = 2x g(x) = 3x³ f(x)/ g(x) = ?

OpenStudy (anonymous):

\[\frac{ 2x }{ 3x^3 }=\frac{ 2 }{ 3 }x^-2=\frac{ 2 }{ 3x^2 }\]

OpenStudy (anonymous):

Because it's a -2 it goes on the bottom?

OpenStudy (anonymous):

because x is a -2*

OpenStudy (anonymous):

yes and for for any base with a negative exponent.. its value its a fraction

OpenStudy (anonymous):

\[x^-1=\frac{ 1 }{ x }\]

OpenStudy (anonymous):

Okay but when it goes on the bottom it becomes a positive again then?

OpenStudy (anonymous):

Yes you are right.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

you are most welcomed